# MDPO: Overcoming the Training-Inference Divide of Masked Diffusion Language Models

### v1

@@ 4 Experiments
    Recognizing Answer Backslide as a manifestation of the training–inference divide, how can we leverage it to close this gap?
@@ 4.1 Experimental Setup
Inference: The performance of MDLMs depends on several inference-time hyperparameter choices. A key distinction lies in whether denoising is performed over the full sequence at once or in blocks. Previous works apply a semi-autoregressive strategy where the sequence is divided into several blocks and generated from left to right. Within each block, tokens are denoised in parallel. We refer to this setting as semi-AR, and contrast it with the setting of denoising all tokens in the entire sequence simultaneously (pure-Diff). In semi-AR, the same remasking strategies used in pure-Diff apply, but each block denoises fewer tokens with fewer steps. To generate a sequence of length $L$ with block size $B$ in $T$ steps, each block is allocated approximately $\lfloor\frac{T\cdot B}{L}\rfloor$ steps. As shown in Fig. 6, performance varies significantly depending on the choice of $B$ and $T$. For controlled comparisons, we adopt two standard configurations throughout our experiments: pure-Diff ($B=L$) and semi-AR ($B=128$). Extensive analysis on choices of block size is in Appendix A.3. Though empirically the pure-Diff setting usually underperforms semi-AR, we keep it as an interesting setting in our main evaluation because of the potential speed advantage that pure-Diff possesses at generating long responses.
Data: We investigate reasoning tasks with answers that can be verified by static reward functions: (1) Mathematical Reasoning where the model generates solutions to mathematical problems and we verify if the ground truth answers are in the solutions with a robust mathematical expression evaluation system (Kydl´ıček & Gandenberger, 2025). Note that different math verifiers can lead to changes in evaluation results and we control the experiments by consistently use the same verifier. (2) Planning with Countdown (Pan et al., 2025), an easy-to-verify combinatorial arithmetic game in which the model tries to reach target numbers using basic arithmetic operations on a given set (3 or 4) of numbers. Evaluation is on MATH-500 (Lightman et al., 2024), a curated dataset of 500 problems from the MATH dataset (Hendrycks et al., 2021); and the Countdown test set by Zhao et al. (2025), specifically. For training on the mathematical reasoning task, we utilize (part of) the math data from Huggingface OpenR1 (Hugging Face, 2025) that consists of 93.7k competition-level math problems from NuminaMath 1.5 (LI et al., 2024). For Countdown, we directly use the respective training split released by Pan et al. (2025).
Model and Training: We compare our methods to two baseline methods: diffu-GRPO (Zhao et al., 2025), which is the first integration of policy gradient methods trained on MDLMs, and LLaDA with supervised finetuning (SFT). We (re)train both the baseline methods and MDPO initialized with LLaDA-8B-Instruct (Nie et al., 2025) under the same settings. Training of all methods is conducted under a fixed compute budget using 8 NVIDIA H100 GPUs with 100 weight update steps. We use a batch size of 128 and apply gradient accumulation when memory constraints occur. For both diffu-GRPO and our proposed MDPO, the group size for group-relative advantage estimation is 8, and the number of denoising steps for sampling rollouts is 128. One special setting that will be discussed in detail in Section 4.3 is that we train MDPO with only Answer Backslide samples, which shows better data efficiency.
@@
+ diffu-GRPO (Zhao et al., 2025) 23.8 34.4 40.2 21.2 26.2 20.6 18.2 35.9 40.1 49.4 55.0 58.9
@@ 4.2 Improvements in Generation Performance and Sample Efficiency
Figure 3: Analysis of samples where Answer Backslide occur. (Left) Maximum length of consecutive correct steps (span) in the denoising trajectory. (Middle) Number of separate correct spans in a trajectory. (Right) Heatmap of the relative position of correct steps across the denoising process.
4.3 Answer Backslide is An Effective Data Filter
We analyse the Answer Backlside phenomenon in more detail in Fig. 3 which reveals two key insights. First, the left and middle subfigures show that correct spans are typically short and fragmented, with trajectories often containing multiple separate correct spans instead of steadily accumulating correct steps. Such fragmentation increases the likelihood of losing correct tokens before reaching the final step. Second, the heatmap shows that correct answers often appear surprisingly very early but tend to decay over subsequent steps instead of steadily accumulating.
These findings not only underline the necessity to address the training-inference divide for MDLMs, but also reveal that Answer Backslide provides highly informative signals of intermediate steps for improving MDLMs. Inspired by this, we train MDPO with only Answer Backslide samples instead of all data so that MDPO can learn from these intermediate signals more efficiently. Specifically, we first use the initialized model to run inference on the whole dataset to identify Answer Backslide samples and then train only on this subset, which constitutes roughly 10% of the original data. The comparison between MDPO-all-data and MDPO in Fig. 4 shows that the model trained with only Answer Backslide excels in most settings, highlighting the effectiveness of Answer Backslide as a data filter for MDPO.
Figure 4: Comparison of MDPO variants. We report final accuracy and proportion of Answer Backslide cases for all models. MDPO-all-data is trained on all data samples whereas MDPO is trained on only Answer Backslide data, both with rollouts sampled from a mixture of semi-AR and pure-Diff. MDPO-pure-Diff and MDPO-semi-AR represent models that are trained on rollouts sampled from only pure-Diff, and semi-AR, respectively.
@@ 4.4 Impact of Sampling Settings on MDPO
The heatmap in  Fig. 3 shows that semi-AR tends to distribute correctness more evenly but with greater fragmentation, while pure-Diff produces an early burst of correctness that decays sharply and rarely recovers. This raises a question of how different sampling settings during the rollout collection of RL affect the final performance. To investigate this, we compare MDPO variants trained on (i) pure-Diff rollouts only, (ii) semi-AR rollouts only, and (iii) an even mixture of both. Fig. 4 shows that training on a single mode yields the largest improvement in that mode’s evaluation setting, but often at the cost of performance in the other mode. The mixture strategy used in our main MDPO setting achieves a more balanced performance, matching or exceeding the best single-mode results in several configurations. This suggests that mixed sampling allows the policy to learn denoising behaviors that generalize across all inference strategies.
5 Conclusion
@@ Acknowledgments
Appendix A Technical Appendices and Supplementary Material
Figure 5: Comparison of different re-masking strategies introduced by MaskGIT.
A.1 Training-Inference Discrepancy
In Section 3.4, we introduce the two remasking strategies from LLaDA, random remasking and low-confidence remasking (LCR). As the tokens are also randomly masked in pre-training, and LCR is usually used as the de-facto standard for inference (Nie et al., 2025), analyzing the difference between random remasking and low-confidence remasking provides us a closer look at the training-inference divide. Table A.3 shows concrete examples to demonstrate the difference between random masking and LCR.
@@ A.1 Training-Inference Discrepancy
A.2 Remasking Scheduling Function
@@
A.3 Block Size in Semi-AR
@@
Table 3: Compare random remasking and confidence-based remasking. We show the model completions at selected steps (16, 32, 48) given a prompt, with the remasked tokens ([M]) by random remasking and LCR with a mask ratio corresponding to the respective step.
@@
[0.2em/0.2em] Masking Ratio 0.75
      Random [M][M][M][M]\[\[M][M][M][M] [M][M][M][M]sum[M][M]
[M][M]1[M]infty \[M][M][M][M][M][M][M][M][M][M] = \[M][M][M][M][M]1[M][M][M][M]{[M]}{k^[M][M][M][M]
$[M][M][M][M]$
[…]
$[M][M][M][M][M][M][M][M]$
$[M][M][M][M][M][M][M][M]$
[M][M][M][M][M][M][M][M] - q}.\]
[…]
[0.2em/0.2em] Masking Ratio 0.5
[…]
$[M][M][M][M][M][M][M][M]$
[M][M] - q}.\]
[…]
[0.2em/0.2em] Masking Ratio 0.25
[…]

### v2

@@ 3 Experiments
    Recognizing over-denoising as a manifestation of the training–inference divide, how can we leverage it to close this gap?
@@ 3.1 Experimental Setup
+ diffu-GRPO (Zhao et al., 2025a) 23.8 34.4 40.2 21.2 26.2 20.6 18.2 35.9 40.1 49.4 55.0 58.9
@@
Inference: The performance of MDLMs depends on several inference-time hyperparameter choices. A key distinction lies in whether denoising is performed over the full sequence at once or in blocks. Previous works apply a semi-autoregressive strategy where the sequence is divided into several blocks and generated from left to right. Within each block, tokens are denoised in parallel. We refer to this setting as semi-AR, and contrast it with the setting of denoising all tokens in the entire sequence simultaneously (pure-Diff). In semi-AR, the same remasking strategies used in pure-Diff apply, but each block denoises fewer tokens with fewer steps. To generate a sequence of length $L$ with block size $B$ in $T$ steps, each block is allocated approximately $\lfloor\frac{T\cdot B}{L}\rfloor$ steps. As shown in Fig. 6, performance varies significantly depending on the choice of $B$ and $T$. For controlled comparisons, we adopt two standard configurations throughout our experiments: pure-Diff ($B=L$) and semi-AR ($B=128$). Extensive analysis on choices of block size is in Appendix B.3. Though empirically the pure-Diff setting usually underperforms semi-AR, we keep it as an interesting setting in our main evaluation because of the potential speed advantage that pure-Diff possesses at generating long responses.
Data: We investigate reasoning tasks with answers that can be verified by static reward functions: (1) Mathematical Reasoning where the model generates solutions to mathematical problems and we verify if the ground truth answers are in the solutions with a robust mathematical expression evaluation system (Kydlíček & Gandenberger, 2025). Note that different math verifiers can lead to changes in evaluation results and we control the experiments by consistently use the same verifier. (2) Planning with Countdown (Pan et al., 2025), an easy-to-verify combinatorial arithmetic game in which the model tries to reach target numbers using basic arithmetic operations on a given set (3 or 4) of numbers. Evaluation is on MATH-500 (Lightman et al., 2024), a curated dataset of 500 problems from the MATH dataset (Hendrycks et al., 2021); and the Countdown test set by Zhao et al. (2025a), specifically. For training on the mathematical reasoning task, we utilize (part of) the math data from Huggingface OpenR1 (Hugging Face, 2025) that consists of 93.7k competition-level math problems from NuminaMath 1.5 (LI et al., 2024). For Countdown, we directly use the respective training split released by Pan et al. (2025).
Model and Training: We compare our methods to two baseline methods: diffu-GRPO (Zhao et al., 2025a), which is the first integration of policy gradient methods trained on MDLMs, and LLaDA with supervised finetuning (SFT). We (re)train both the baseline methods and MDPO initialized with LLaDA-8B-Instruct (Nie et al., 2025) under the same settings. Training of all methods is conducted under a fixed compute budget using 8 NVIDIA H100 GPUs with 100 weight update steps. We use a batch size of 128 and apply gradient accumulation when memory constraints occur. For both diffu-GRPO and our proposed MDPO, the group size for group-relative advantage estimation is 8, and the number of denoising steps for sampling rollouts is 128. One special setting that will be discussed in detail in Section 3.3 is that we train MDPO with only over-denoising samples, which shows better sample efficiency.
@@ 3.2 Improvements in Generation Performance and Sample Efficiency
3.3 Over-denoising as An Effective Data Filter
Figure 3: Analysis on the occurrence of over-denoising samples across multiple inference settings (Block size, Step). (Left) Maximum length of consecutive correct steps (span) in the denoising trajectories. (Middle) Number of separate correct spans over trajectories. (Right) Heatmap of the relative position of correct steps across the denoising process.
We analyze the over-denoising phenomenon in more detail in Fig. 3 which reveals two key insights. First, the left and middle subfigures show that correct spans are typically short and fragmented, with trajectories often containing multiple separate correct spans instead of steadily accumulating correct steps. Such fragmentation increases the likelihood of losing correct tokens before reaching the final step. Second, the heatmap shows that correct answers often appear surprisingly very early but tend to decay over subsequent steps instead of steadily accumulating.
These findings not only underline the necessity to address the training-inference divide for MDLMs, but also reveal that over-denoising provides highly informative signals of intermediate steps for improving MDLMs. Inspired by this, we train MDPO with only over-denoising samples instead of all data so that MDPO can learn from these intermediate signals more efficiently. Specifically, we first use the initialized model to run inference on the whole dataset to identify over-denoising samples and then train only on this subset, which constitutes roughly 10% of the original data. The comparison between MDPO-all-data and MDPO in Fig. 4 shows that the model trained on only over-denoising excels in most settings, highlighting the effectiveness of over-denoising as a data filter for MDPO.
@@ 3.4 Impact of Sampling Settings on MDPO
Figure 4: Comparison of MDPO variants. We report the final accuracy and proportion of over-denoising cases for all models. MDPO-all-data is trained on all data samples, whereas MDPO is trained on only over-denoising data, both with rollouts sampled from a mixture of semi-AR and pure-Diff. MDPO-pure-Diff and MDPO-semi-AR represent models that are trained on rollouts sampled from only pure-Diff, and semi-AR, respectively.
The heatmap in  Fig. 3 reveals that semi-AR tends to distribute correctness more evenly but with greater fragmentation, while pure-Diff produces an early burst of correctness that decays sharply and rarely recovers. This raises a question of how different sampling settings during the rollout collection of RL affect the final performance. To investigate this, we compare MDPO variants trained on (i) pure-Diff rollouts only, (ii) semi-AR rollouts only, and (iii) an even mixture of both. Fig. 4 shows that training on a single mode yields the largest improvement in the evaluation setting of the respective mode, but often at the cost of performance in the other modes. The mixture strategy used in our main MDPO setting achieves a more balanced performance, matching or exceeding the best single-mode results in several configurations. This suggests that mixed sampling allows the policy to learn denoising behaviors that generalize across all inference strategies.
4 Related Work
Diffusion probabilistic models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have become the de facto standard for generative modeling of continuous signals, including images, videos, 3D shapes, and robotic trajectories. Building on early efforts in discrete diffusion (Austin et al., 2021; Hoogeboom et al., 2021), several works have adapted diffusion models for text generation. Diffusion-LM (Li et al., 2022) and Self-Conditioned Embedding Diffusion (Strudel et al., 2022) have pioneered embedding-based diffusion approaches. Liu et al. (2025) introduces a novel framework that separates the generation process into two models: a planner and a denoiser. Meanwhile, masked diffusion (Austin et al., 2021; Sahoo et al., 2024; Ou et al., 2025) has been established as a special case of discrete diffusion, with recent research significantly scaling their capabilities. Notably, DiffuLLaMA (Gong et al., 2025) initializes masked diffusion language models using pre-trained LLaMA weights. LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025) scale MDLMs by pre-training from scratch to the billion-parameter regime, achieving performance comparable to similarly sized AR LLMs.
Training diffusion models with RL has been primarily explored in image diffusion models (Black et al., 2024; Fan et al., 2023) on non-differentiable objectives such as human-perceived image quality. Represented by Group Relative Policy Optimization (GRPO) (Shao et al., 2024), RL has emerged as a promising tool to optimize sparse and task-specific reward signals for language models (Guo et al., 2025; Yu et al., 2025b; Luo et al., 2025b; a; Hu et al., 2025; Zeng et al., 2025), especially for post-training. These RL algorithms assume sequential token-by-token generation, a key property of AR models. However, MDLMs differ fundamentally: they denoise entire sequences from corrupted inputs, making AR-specific RL inapplicable. Recent attempts Zhao et al. (2025a; b) adapt GRPO to MDLMs with inserted partial ground-truth reasoning traces to improve sample efficiency, meanwhile concurrent work Wang et al. (2025b) also observes the over-denoising phenomenon and adds entropy-based rewards to optimize MDLMs with RL. In contrast, we introduce MDPO, a policy optimization method designed for MDLMs to learn effective denoising trajectories by optimizing the iterative denoising process as a multi-step decision-making problem. To our knowledge, MDPO is the first RL method explicitly designed to optimize MDLMs denoising trajectories as multi-step decision-making.
Kim et al. (2025); Wang et al. (2025a) show that though MDLMs face inherently harder training challenges due to solving computationally intractable infilling tasks, using adaptive decoding strategies effectively alleviates these. Following this, Block Diffusion (Arriola et al., 2025) models sequences in autoregressive blocks with intra-block diffusion, concurrent work Li et al. (2025) observes the same over-denoising phenomenon and applies a confidence-based early-commit strategy for faster denoising. Our proposed Running Confidence Remasking enables more flexible decoding to substantially enhance the denoising quality by enabling iterative revision of early-predicted ‘noisy’ tokens.
5 Discussion & Conclusion
@@ Acknowledgments
Appendix
Appendix A Experimental Settings
A.1 Advantage Estimation
In Eq. (6), we define the score of each generation as
$s_{t,k}=r(x_{t,k})-r(x_{t+1,k})+\frac{1}{t}\sum_{j=0}^{t-1}r(x_{j,k})$ (11)
to compute the advantage. The intuition behind this is the first part of the definition $r(x_{t,k})-r(x_{t+1,k})$ incentivizes immediate improvement over the previous, noisier prediction and the second part $\frac{1}{t}\sum_{j=0}^{t-1}r(x_{j,k})$ encourages predictions that steer the trajectory toward cleaner states that score well overall. We highlight that the design of advantage estimation plays a pivotal role in determining the final performance, since poorly structured estimators can be easily hacked, allowing the model to overfit to superficial reward patterns rather than acquiring the intended behavior.
A.2 KL Divergence Approximation
We approximate the KL divergence between the trained model and the reference model $\mathbb{D}_{KL}\left[p_{\theta}||p_{\text{ref}}\right]$ in Eq. (4) following Schulman (2020). Define $d=\frac{p_{\text{ref}}(x)}{p_{\theta}(x)}$, three estimators are defined in Schulman (2020):
$k1=-\log d\qquad k2=\frac{(\log d)^{2}}{2}\qquad k_{3}=(d-1)-\log d$ (12)
where $k1$ is an unbiased but high-variance estimator, as it is negative for half of the samples, while the KL divergence should always be positive. Using such an estimator inside Eq. (4) where the KL divergence is deducted from the clipped advantage may artificially reward models for ‘negative divergences’, producing biased optimization dynamics. Therefore, we seek an estimator that is always non-negative, for robustness and theoretical consistency with the definition of KL. $k2$ and $k3$ are both non-negative, and $k2$ is more biased than $k3$, according to Schulman (2020). In our preliminary experiments, we found small differences in the performance between using $k2$ and $k3$. Therefore, we choose $k2$ as the final KL approximation in our current implementation for the simplicity of computation.
A.3 Hyperparameter Setting & Other Technical Details
Here we list other technical details that are necessary to reproduce our experimental results. We set the clipping parameter $\epsilon=0.2$ in the objective in Eq. (5) and the KL divergence scaling factor $\beta=0.08$ in Eq. (4). In our experiments, we observe that a larger $\beta$ prevents the model from overfitting to the training data, given that we are supervising the intermediate trajectories and the model receives multiple supervisions for each data sample. Another implementation choice we make to avoid overfitting is to reduce the number of intermediate denoising steps used to create the gradients. Despite the expectation is over all the denoising steps as defined in Eq. (4), we sort the denoising steps according to their absolute advantage values and pick the largest $c$ steps. This not only prevents the model from overfitting by reducing the number of the same data samples are trained on, but also reduces the computation time. $c$ is usually set to $8$ in our experiments. The intuition behind sorting advantages with absolute values is to supervise the model with both good and bad denoising. Learning rate is set to 7e-7.
In Section 3.3, we find that using only over-denoising samples is more data-efficient and produces better results. Although this leads to overhead in pre-computing the over-denoising samples before training, it is still computationally efficient in practice. The reason is that for each inference setting, we only need to infer over the whole training set once and use the over-denoising data samples repetitively for sampling rollouts in the online RL in one training or over multiple training sessions. As the computation cost of inference is much less than completing forward and backward propagation in training, the overhead created by inference is negligible compared to the computation cost for training on all data.
Appendix B Model Analysis
Figure 5: Comparison of different re-masking strategies introduced by MaskGIT. The performance of LLaDA with changing re-masking strategies is not consistent with MaskGIT, a image generation model.
B.1 Training-Inference Discrepancy
In Section 2.4, we introduce the two remasking strategies from LLaDA, random remasking and low-confidence remasking (LCR). As the tokens are also randomly masked in pre-training, and LCR is usually used as the de-facto standard for inference (Nie et al., 2025), analyzing the difference between random remasking and low-confidence remasking provides us a closer look at the training-inference divide. Table B.4 shows concrete examples to demonstrate the difference between random masking and LCR.
@@ B.1 Training-Inference Discrepancy
B.2 Remasking Scheduling Function
@@
B.3 Block Size in Semi-AR
@@
Figure 7: Comparison of LLaDA and Dream on MATH. Results indicate that the performance of Dream is highly dependent on the number of denoising steps, while LLaDA benefits more from the semi-AR setting (with block length 128).
B.4 LLaDA and Dream
As an extended analysis, we choose another pre-trained MDLM Dream 7B (Ye et al., 2025) to compare with LLaDA. Note that Dream 7B builds on a different algorithm to LLaDA and initializes from an auto-regressive (AR) checkpoint Qwen2.5 7B (Yang et al., 2024). Though the algorithm still follows the absorbing discrete diffusion framework to transit tokens to and from the predefined absorbing state (masking token), Dream differs in predicting the next token and filling the next token with the prediction if the next token is masked. This technical difference originates from the fact that Dream is adapted from an AR model which is trained on next token prediction. We observe a clear behavioral difference between the pre-trained from scratch LLaDA and AR-adapted Dream, on MATH task shown in Fig. 7.
The first observation is that the performance of Dream is highly dependent on the denoising steps, but not the block length. In the semi-AR setting (Block Length of 128 in Fig. 7), which substantially improves the performance of LLaDA, Dream achieves downstream performance similar to the pure-Diff setting. We hypothesize that this arises because the model is adapted from an AR model, where earlier tokens consistently receive higher confidence. Consequently, pure-Diff effectively converges toward semi-AR, since tokens in early positions typically obtain higher confidence scores and are rarely remasked. Evidence for this convergence is shown in Fig. 8, where Dream predominantly assigns higher confidence to earlier masked tokens.
Our second observation is that Dream exhibits a smaller degree of over-denoising compared to LLaDA. We hypothesize that this difference also stems from its adaptation from AR models. While we see over-denoising as a symptom of the training–inference divide inherent in masked diffusion, Dream is pre-trained on a mixture of objectives, including both next-token prediction and masked diffusion. In particular, the pre-training of Dream on the masked diffusion objective consumes 580 billion tokens, whereas the original initialization Qwen 2.5 pre-training covers 18 trillion tokens. We argue that this imbalance in the scale of pre-training tokens between the two stages reduces the extent to which Dream experiences the training–inference divide, making its behavior more similar to that of an AR model. As a result, the discrepancy is less pronounced in Dream, and applying our proposed methods might yield only marginal improvements.
Figure 8: Quantitative analysis of predicted and ‘kept’ (not remasked) tokens in LLaDA and Dream. We apply confidence-based remasking with pure-Diff to both models and plot the distribution of retained tokens by their rank among masked positions at respective denoising steps across trajectories. The results show that Dream tends to assign higher confidence to earlier masked tokens, resembling AR models, whereas LLaDA exhibits a two-sided pattern, giving higher confidence to both early and late masked tokens.
Table 3: Comparison of random remasking and confidence-based remasking. We show the model completions at selected denoising steps (16, 32, 48) given a prompt, with the remasked tokens ([M]) by random remasking and LCR with a mask ratio corresponding to the respective step.